### v1

# LC-LLM: Explainable Lane-Change Intention and Trajectory Predictions with Large Language Models

Abstract
To ensure safe driving in dynamic environments, autonomous vehicles should possess the capability to accurately predict the lane change intentions of surrounding vehicles in advance and forecast their future trajectories. Existing motion prediction approaches have ample room for improvement, particularly in terms of long-term prediction accuracy and interpretability. In this paper, we address these challenges by proposing LC-LLM, an explainable lane change prediction model that leverages the strong reasoning capabilities and self-explanation abilities of Large Language Models (LLMs). Essentially, we reformulate the lane change prediction task as a language modeling problem, processing heterogeneous driving scenario information in natural language as prompts for input into the LLM and employing a supervised fine-tuning technique to tailor the LLM specifically for our lane change prediction task. This allows us to utilize the LLM’s powerful common sense reasoning abilities to understand complex interactive information, thereby improving the accuracy of long-term predictions. Furthermore, we incorporate explanatory requirements into the prompts in the inference stage. Therefore, our LC-LLM model not only can predict lane change intentions and trajectories but also provides explanations for its predictions, enhancing the interpretability. Extensive experiments on the large-scale highD dataset demonstrate the superior performance and interpretability of our LC-LLM in lane change prediction task. Compared with baseline models, the performance of intention prediction can be improved by 3.1%, the root-mean-squared error (RMSE) of lateral trajectory prediction can be reduced by 19.4%, and the RMSE of longitudinal trajectory prediction can be reduced by 38.1%. To the best of our knowledge, this is the first attempt to utilize LLMs for predicting lane change behavior. Our study shows that LLMs can encode comprehensive interaction information for driving behavior understanding.
Index Terms:
Intention prediction, Trajectory prediction, Lane change, Fine-tuning, Interpretability, Large language models, Autonomous driving.
I Introduction
Lane change prediction is a critical task for autonomous driving systems to ensure safe and efficient navigation on the road. Lane changing behavior, which refers to the action of a vehicle moving from one lane to another on the road, is one of the most complex driving behaviors due to intricate interactions with other vehicles. Accurately forecasting the underlying lane change intentions and future trajectories of surrounding vehicles in advance is essential for enabling autonomous vehicles to make informed decisions and take appropriate actions to avoid collisions and maintain smooth traffic flow.
In the past five years, considerable advancements have been made in the field of lane change prediction. Various approaches have been explored to enhance performance in detecting lane change maneuvers [1, 2, 3]. However, at the early stage of a lane change maneuver when lateral displacement is minimal, it is essential to rely on the interaction information between the target vehicle and its surrounding vehicles for accurate predictions. Therefore, the ability to effectively model interaction information among vehicles becomes paramount. Previous methods exhibit limited capability in understanding interactive information, which results in weak long-term prediction performance. Additionally, lane change trajectory prediction has witnessed notable progress with the adoption of deep learning techniques, such as Graph Neural Networks (GNN) and Transformers, leading to competitive results [3, 4, 5]. Despite these successes, these deep learning-based approaches frequently suffer from a lack of interpretability, as they generate predictions about future behaviors without offering substantial explanations for their results. In summary, enhancing the capability to model interactions among vehicles and augmenting the interpretability of prediction results are two significant challenges in the domain of lane change prediction for autonomous driving systems.
In recent years, advancements in the field of Large Language Models (LLMs) have opened new avenues for addressing challenges associated with lane change prediction. Noteworthy progress in LLMs [6, 7, 8] has showcased their robust information comprehension skills and powerful common sense reasoning abilities. Subsequent research by Mao et al. [9], Chen et al. [10] and Wu et al. [11] has successfully leveraged these LLMs to confront challenges in motion planning and perception, highlighting the considerable potential of LLMs in the domain of autonomous driving. However, current researches on lane change prediction have not yet incorporated LLMs. Given the LLMs’ profound understanding of complex driving scenarios and their extensive knowledge base, including common driving knowledge, it is rational to utilize LLMs in modeling interactions between vehicles for the enhancement of future driving behavior predictions. In addition, LLMs have demonstrated the capacity to leverage their vast knowledge base to generate explanations for their predictions [12, 13]. Inspired by these studies, there is potential to leverage LLMs for the explanation of predicted lane change intentions and future trajectories, thereby augmenting the interpretability of predictions in autonomous driving systems.
In this paper, we propose a novel approach that leverages LLMs’ powerful reasoning capabilities and self-explanation capabilities to address lane change prediction challenges mentioned above. In essence, we reformulate the task of predicting lane change intentions and trajectories as a language modeling problem. To this end, we integrate heterogeneous inputs, including target vehicle state, surrounding vehicle state, map information, and Chain-of-Thought (CoT) statements, converting them into structured prompts in natural language to input into the LLM. Then, we employ a supervised fine-tuning technique to tailor the LLM specifically for the lane change prediction task. Furthermore, we integrate explanatory requirements into the prompts in the inference stage, thus enabling our fine-tuned model to generate explanations for the lane change prediction results. Benefiting from the powerful capabilities of LLM, our fine-tuned model demonstrates better performance and interpretability in lane change prediction.
The main contributions of this paper include:
  • We propose LC-LLM, the first Large Language Model for lane change prediction. It leverages the powerful capabilities of LLMs to understand complex interactive scenarios, enhancing the performance of lane change prediction.
  • Our LC-LLM achieves explainable predictions. It not only predicts lane change intentions and trajectories but also generates explanations for the prediction results.
  • Extensive experiments on the highD dataset demonstrate that our LC-LLM outperforms all baseline models, showcasing superior performance and interpretability in lane change prediction.
The remainder of the paper is organized as follows. Section II reviews previous research related to our work. Section III presents the LLM-based lane change intentions and trajectory prediction approach. Section IV provides experimental results demonstrating the performance of the proposed approach. Finally, the conclusions are drawn in Section V.
Figure 1: The pipeline of our LC-LLM. The green trajectory in the scenario on the left is the future trajectory of the target vehicle, which the model aims to predict. Observations are described using natural language prompts, which are input into the LLM. Supervised fine-tuning methods are then applied to fine-tune the LLM for accurate prediction of lane change intentions and trajectories. During the inference phase, prompts are designed to include explanatory requirements. As a result, the fine-tuned model is capable of predicting the target vehicle’s lane change intentions and future trajectory in the current frame, while simultaneously providing explanations for its predictions, thus improving the interpretability of the model’s outputs.
Figure 2: A lane change scenario. The target vehicle is depicted in green and the surrounding vehicles are depicted in blue. The orange line illustrates the trajectory of the target vehicle in the future $t$ timesteps. The advanced prediction time T denotes the temporal interval between the lane change frame and the current frame.
Figure 3: An example of our input prompts in the fine-tuning stage. The input prompts comprise a system message displayed in the upper text block and a user message presented in the lower text block. The user message mainly contains map information, the target vehicle’s state, the spatial relationships between the target vehicle and its surrounding vehicles, as well as constructed chain-of-thought statements.
II RELATED WORK
The literature reviews provided by Huang et al. [14] and Mozaffari et al. [15] discuss various methods for predicting the future states of vehicles for autonomous driving systems. These prediction studies can be classified into the following two categories based on the output type: Intention prediction and trajectory prediction. Concurrently, LLMs have shown rapid progress in recent developments. Comprehensive overviews of LLMs applications in autonomous driving are discussed in [16].
II-A Intention Prediction
Lane change intention prediction is a critical component in the domain of autonomous driving and advanced driver assistance systems (ADAS). Research in this field focuses on accurately predicting vehicles’ intention to change lanes before the maneuver is executed, thereby enhancing road safety and traffic flow. Various approaches have been explored, including machine learning techniques, probabilistic modeling, and deep learning. Mandalia et al. [17] first applied the support vector machine (SVM) to identify lane change maneuvers by incorporating features such as acceleration, steering angle, distance, and so on. Lyu et al. [18] predict lane change intentions based on the support vector machine-recursive feature elimination (SVM-RFE) model. He et al. [1] designed a Dynamic Bayesian Network (DBN) with the purpose of distinguishing between vehicle-following and lane-change maneuvers. Further advancements were made by Hong et al. [19] and Mozaffari et al. [2], who utilized a convolutional neural network (CNN) as feature extractors to fuse complex driving environment context and effectively learn the driving behavior. Zyner et al. [20, 21] employed Long Short-Term Memory Networks (LSTM) and Recurrent Neural Networks (RNN), respectively, as a sequence classifier for predicting vehicles’ driving intentions. Xin et al. [22] implemented a dual-block LSTM architecture, where the first LSTM block processes sequential trajectory data to recognize driver intentions as an intermediate indicator. Izquierdo et al. [23] utilized a hybrid CNN-LSTM model to capture both local and global contextual features, as well as temporal information, to forecast vehicle lane change intentions. Gao et al. [3] proposed a dual Transformer model, which includes a lane change intention prediction model and a trajectory prediction model. Although these approaches have been proven to have good performance in detecting lane change maneuvers that have already started, the capability of predicting driving intention in advance needs to be improved.
II-B Trajectory Prediction
Research on trajectory prediction for autonomous driving is a critical area of study, aiming to enhance the safety and efficiency of self-driving vehicles by forecasting the future trajectory of surrounding entities such as cars, cyclists, and pedestrians. Early approaches used bird’s-eye view images as the input and applied a CNN framework to process the rasterized scene for trajectory prediction [24, 25]. Furthermore, several studies employed LSTM networks, utilizing one LSTM as an encoder to capture features from historical trajectories and another LSTM as a decoder to predict future trajectories [26, 27, 22]. More recent works represented scenes with vectorized data such as points and polylines, and processed them with GNNs [28, 29] or Transformers [4, 3, 30] to effectively model the interactions between traffic participants and environment. In [29], Gao et al. proposed VectorNet which used GNNs to extract features from vectorized high-definition (HD) maps, thereby avoiding lossy rendering and computationally intensive CNN encoding. In [4], Shi et al. proposed the Motion transformer framework, which models motion prediction as the joint optimization of global intention localization and local movement refinement. A recent study [5] represented continuous trajectories as sequences of discrete motion tokens, framing multi-agent motion prediction as a language modeling task. While these deep learning-based approaches achieve competitive results, their predictions often lack interpretability, which is not conducive to the development of safer and more transparent autonomous driving systems.
II-C LLMs for Autonomous Driving
Recent advancements in LLMs have demonstrated remarkable capabilities in extensive knowledge storage, logical reasoning, and question-answering. Considering these competencies, it is a natural extension to contemplate the application of LLMs to enhance the field of autonomous driving [16]. Wu et al. [11] integrated LLMs with 3D detection and tracking tasks by using language prompts as semantic cues. Mao et al. [9] proposed a prompting-reasoning-finetuning strategy, enabling LLM to generate driving trajectories and showcased the detailed numerical reasoning abilities of GPT-3.5 in motion planning. Chen et al. [10] designed an object-level multimodal LLM architecture that combines vectorized numeric modalities with a pre-trained LLM. Additionally, Xu et al. [31] presented an interpretable autonomous driving system employing LLMs and developed a visual instruction tuning dataset for interpretable autonomous driving. While LLMs have been extensively explored in the field of autonomous driving perception and planning, their application in the domains of intention prediction and trajectory prediction remains unexplored.
III Methodology
In this section, we introduce our LC-LLM, a model based on LLM designed for predicting lane change intentions and trajectory in autonomous driving systems. The whole pipeline of our LC-LLM is depicted clearly in Fig. 1. We reconceptualize the task of predicting intentions and trajectory as a language modeling problem. To this end, we articulate observations using natural language as prompts for input into the LLM and leverage supervised fine-tuning techniques to tailor the LLM to this specific task. During the inference stage, we incorporate explanatory requirements into the prompt. As a result, our fine-tuned model, LC-LLM, not only forecasts the lane change intentions and future trajectories but also provides explanations for the predictions, thereby enhancing their interpretability.
III-A Problem Formulation
In this paper, we aim to develop a predictive model based on LLM that is designed to simultaneously predict the lane change intention and the trajectory of the target vehicle. A representative lane change scenario is depicted in Fig. 2, where the target vehicle is illustrated in green and the surrounding vehicles in blue. The objective is to forecast the position of the target vehicle’s trajectory at t future timesteps, as well as to determine its lane change intentions within the same temporal window, as indicated by the orange line in Fig. 2.
We define the state of the target vehicle as tv, the states of the surrounding vehicles as sv, and the map information as m, which serve as the inputs to our model. The outputs are the lane change intention $\mathcal{I}$ and the future trajectory $\mathcal{T}$. The process of the model $\mathcal{F}$ can be formulated as follows:
$\mathcal{I},\mathcal{T}=\mathcal{F}(\textbf{{tv}},\textbf{{sv}},\textbf{{m}})$ (1)
The $\mathcal{I}\in\{0,1,2\}$, where 0 denotes “keep lane”, 1 denotes ”left lane change”, and 2 denotes ”right lane change”. The future trajectory of prediction horizon $t$ can be denoted by $\mathcal{T}=\{(\textit{x}_{1},\textit{y}_{1}),(\textit{x}_{2},\textit{y}_{2}),...,(\textit{x}_{t},\textit{y}_{t})\}$. The target vehicle state tv consists of a historical trajectory of this vehicle and its current velocity and type. The surrounding vehicle states m contain each vehicle’s current velocity, type, and their spatial relationship relative to the target vehicle. The map information m includes details such as lane identifiers and lane markings.
In this paper, we reframe the task of predicting intentions and trajectory as a language modeling problem. By utilizing natural language to describe both the input data and prediction results, we can represent them as a sequence of tokens. The input data which describe the current frame driving scenario can be denoted as a sequence of tokens $\textit{T}_{s}$:
$\textit{T}_{s}=K(\textbf{{tv}},\textbf{{sv}},\textbf{{m}})$ (2)
The prediction result which describes the lane change intentions and trajectory can also be represented as a sequence of tokens $\{\textit{T}_{1},\textit{T}_{2},...,\textit{T}_{n}\}$:
$\{\textit{T}_{1},\textit{T}_{2},...,\textit{T}_{n}\}=K(\{\mathcal{I},(\textit{x}_{1},\textit{y}_{1}),(\textit{x}_{2},\textit{y}_{2}),...,(\textit{x}_{t},\textit{y}_{t})\})$ (3)
where the $K$ is a language tokenizer used to transform input data and output results into tokens. The $\textit{T}_{i}$ represents the $i$-th token in the sequence. By adopting this language-based representation, we can redefine the prediction problem as a language modeling problem, with the loss function resembling that of language modeling [32]:
$\mathcal{L}=-\sum_{i=1}^{n}\log\textit{P}(T_{i}^{*}|T_{s},T_{<i})$ (4)
where $\textit{T}_{i}^{*}$ denotes the ground truth of the next token given its historical tokens, and $n$ represents the tokens number of ground truth in one sample. By maximizing the conditional probability P of the token $\textit{T}_{i}^{*}$, our model can effectively predict lane change intentions and future trajectory.
III-B Prompting
Generally, LLMs receive inputs in the form of natural language rather than unprocessed vectorized data. Consequently, the formulation of effective prompts that describe the current observations in natural language is critical. Several studies have sought to harness the deep reasoning capabilities of the LLMs through clever prompt design [9, 33]. Building on these prior efforts, we have crafted prompts that are clearer, more intelligent, and better structured. Fig. 3 provides an example of our input prompts.
As illustrated in Fig. 3, the input prompts consist of a system message displayed in the upper text block and a user message presented in the lower text block. The system message maintains consistency across diverse driving scenarios. It delineates the designated role of the LLM, provides details of the coordinate system, and outlines the information and format for the LLM’s output. In our paper, the assigned role for the LLM is that of a predictive model integrated within an autonomous driving system. The coordinate system in each driving scenario is the vehicle coordinate system, which is centered on the current position of the target vehicle. The expected output includes predictions of lane change intentions and trajectory points over a future time horizon of four seconds.
The user message provides a description of the observations specific to the current frame, thus varying with each driving scenario. It includes information about the map, the state of the target vehicle, the spatial relationships between the target vehicle and its surrounding vehicles, as well as constructed Chain-of-Thought (CoT) [34] statements. Map information in the prompts primarily denotes the number of lanes in the scenario and indicates whether the target vehicle is located in the leftmost, middle, or rightmost lane. Prompts related to the target vehicle’s state are generated by detailing the target vehicle’s historical trajectory over the past two seconds, its current velocity, and its vehicle type. Given that most vehicles do not have a large lateral displacement 4 seconds before changing lanes, the model mainly relies on the interaction information between the target vehicle and surrounding vehicles to predict lane change intentions in advance. Therefore, comprehending the information pertaining to surrounding vehicles is crucial for accurately predicting the lane change intentions of the target vehicle. We denote the information of the nearest vehicles in eight directions surrounding the target vehicle as surrounding vehicle prompts. These directions include the front, left front, right front, left side, right side, rear, left rear, and right rear. The surrounding vehicle information in each direction encompasses details such as vehicle type, current speed, and relative distance from the target vehicle. This comprehensive surrounding vehicle information serves as a critical prompt for the predictive model, enabling a thorough analysis of the contextual interactions influencing the target vehicle’s lane change intentions.
Recently, CoT prompting has demonstrated remarkable capabilities in carrying out more intricate reasoning tasks [35, 34, 36]. Drawing inspiration from Zero-shot-CoT [34] and Least-to-Most prompting, we formulate CoT prompts by subdividing the task into three subproblems. These subproblems serve as guidance for the LLM in the reasoning process, facilitating the subsequent resolution of the final prediction problem. In the aforementioned prompt, a wealth of information is provided to the LLM. However, the LLM may not thoroughly comprehend this information or may resort to simplistic problem-solving strategies. Therefore, the incorporation of CoT prompts is deemed necessary to instruct the LLM to deliberate thoughtfully on aspects such as target vehicle information, potential intentions of surrounding vehicles, and the spatial relationships and interactions between the target vehicle and its surroundings. The details of CoT prompts are shown in Fig. 3.
The last portion of our prompts details the content and format of the final answer. Our model’s objective is to accurately forecast lane change intentions and trajectories. Consequently, the final answer should encapsulate two tasks: intention prediction and trajectory prediction. It is noted in the literature that high-level intention prediction substantially contributes to trajectory prediction [22, 3]. Therefore, aligning with the characteristics of the autoregressive language model, we have formulated the final answer to sequentially output the intention first, followed by the corresponding future trajectory points.
Figure 4: A data sample of Llama format. Special tokens, such as [INST], serve as delimiters distinguishing system messages, user messages, and answer messages. The details of these messages are elucidated in the Prompting section.
Figure 5: An example of our input prompts in the inference stage. The content in the green text block is the added prompts in the inference stage. The ellipsis part refers to other content of the input prompts, the same as Fig. 3.
III-C Fine-tuning
In our work, we utilize an open-source foundational language model, Llama-2-7b-chat [8], as our pre-trained LLM. To achieve parameter-efficient finetuning, we adopt the Low-Rank Adaptation (LoRA) [37] strategy, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture. Additionally, we customize the LLM for our specific prediction task using supervised fine-tuning techniques. The raw dataset utilized in our work originates from the highD dataset [38], which captures human naturalistic vehicle trajectories on German highways. Following the fine-tuning instructions provided for Llama-2, each data sample is formatted to include an input prompt and a corresponding answer, separated by a special token. This formatting is illustrated in the Fig. 4. The answer for each sample is derived from the ground truth obtained from the highD dataset, encompassing future trajectories and intentions for programmatic automatic labeling. In summary, we process highD raw dataset with natural language, format each sample into the Llama format, and then feed it into LLM. Finally, we fine-tune the LLM by aligning the LLM’s output $\{\mathcal{I},\mathcal{T}\}$ with the corresponding ground truth labels $\{\mathcal{I}^{*},\mathcal{T}^{*}\}$. This fine-tuning process is executed through the language modeling loss $\mathcal{L}$ as defined in Equation 4. During fine-tuning, we mask the loss for tokens within the input prompts, focusing backpropagation only on the tokens comprising the answers, in alignment with the fine-tuning approach of Llama-2. By following this approach, our fine-tuned model, LC-LLM, can predict human driving behaviors, including lane change intentions and future trajectories. This capability is crucial for ensuring the safety of ego vehicles within an autonomous driving system.
III-D Interpretability of Prediction
A prevalent limitation in contemporary autonomous driving prediction models lies in their constrained interpretability. This is attributed to the fact that these models generate predictions about the future behaviors or trajectories of the target vehicle through black-box neural networks, offering little explanation behind their prediction results. In our research, we address this limitation by incorporating explanatory requirements into the input prompts in the inference stage, where the weights of our fine-tuned model are fixed. An example of these input prompts during the inference stage is illustrated in Fig. 5. Benefiting from the self-explanation [12, 13] capabilities of LLMs, our fine-tuned model, LC-LLM, predicts lane change intentions and future trajectories and also provides explanations for its prediction results in the inference stage, thereby enhancing the interpretability of the prediction results. This approach facilitates a more transparent and understandable prediction process, which is crucial for the practical application of autonomous driving systems. Some visual examples of the interpretability of our LC-LLM model are shown in Fig. 8.
IV Experimental Results and Analysis
This section outlines the evaluations of our proposed LC-LLM model. First, we detail the dataset processing and the experimental setup. Subsequently, we elucidate the evaluation metrics. Next, we compare the proposed model with the baseline models and analyze the results quantitatively. Moreover, we perform an ablation study on the key components of our proposed method. Finally, we conduct an interpretability study and a zero-shot study to evaluate the interpretability and effectiveness of our approach.
TABLE I: Hyperparameters for training our LC-LLM.
Hyper-parameter Value
Learning rate 5e-4
Batch size 8
Train epoch 2
Lora_r 64
Lora_alpha 16
Gradient_accumulation_steps 8
Load_in_bits 4
TABLE II: Comparison of the intention prediction performance of the proposed LC-LLM model and baseline models on the highD dataset.
Model Intention $T\in[0,1]$ $T\in(1,2]$ $T\in(2,3]$ $T\in(3,4]$ Avg. ($T\in[0,4]$)
P(%) R(%) F1(%) P(%) R(%) F1(%) P(%) R(%) F1(%) P(%) R(%) F1(%) P(%) R(%) F1(%)
LSTM[22] LK 84.6 97.0 90.4 80.5 97.7 88.3 69.6 97.5 81.2 56.9 97.4 71.8 71.2 97.4 82.3
LLC 97.8 83.0 89.8 98.6 79.3 87.9 97.1 67.3 79.5 95.3 49.9 65.5 97.4 69.9 81.4
RLC 98.9 99.4 99.1 98.7 97.0 97.8 98.8 89.6 94.0 97.6 74.7 84.6 98.5 90.2 94.2
Macro avg. 93.8 93.1 93.1 92.6 91.3 91.3 88.5 84.8 84.9 83.3 74.0 74.0 89.0 85.8 85.9
Transformer[3] LK 89.9 91.7 90.8 85.5 91.2 88.4 74.3 91.1 81.8 63.4 91.6 75.0 76.9 91.4 83.5
LLC 95.6 91.1 93.3 94.6 85.7 89.9 93.5 73.8 82.4 92.7 62.4 74.6 94.2 78.2 85.5
RLC 96.0 98.5 97.2 95.4 98.3 96.8 95.2 93.7 94.4 93.2 82.2 87.3 95.0 93.1 94.1
Macro avg. 93.8 93.8 93.8 91.9 91.7 91.7 87.6 86.2 86.2 83.1 78.7 79.0 88.7 87.6 87.7
LC-LLM-7B LK 98.0 81.5 89.0 97.7 81.7 88.9 86.7 80.3 83.4 70.8 80.4 75.3 86.8 81.0 83.8
LLC 90.7 98.0 94.2 90.5 97.5 93.9 87.2 89.6 88.4 85.1 75.8 80.2 88.6 90.2 89.2
RLC 91.9 100.0 95.8 91.6 99.7 95.5 91.0 95.1 93.0 86.6 84.4 85.4 90.4 94.8 92.5
Macro avg. 93.5 93.2 93.0 93.3 92.9 92.8 88.3 88.4 88.3 80.8 80.2 80.3 88.6 88.7 88.6
LC-LLM-13B LK 98.2 89.1 93.5 96.3 88.6 92.3 86.8 88.7 87.8 68.7 90.3 78.0 85.7 89.2 87.4
LLC 91.9 98.6 95.1 91.2 96.7 93.9 89.4 89.7 89.5 86.3 75.7 80.7 89.9 90.2 90.0
RLC 97.6 99.7 98.6 96.9 98.8 97.8 96.5 94.1 95.3 95.2 76.9 85.1 96.6 92.3 94.4
Macro avg. 95.9 95.8 95.7 94.8 94.7 94.6 90.9 90.8 90.9 83.4 81.0 81.3 90.7 90.6 90.6
LC-LLM-70B LK 96.4 92.5 94.4 96.1 91.8 93.9 78.9 92.5 85.2 64.0 93.4 76.0 81.4 92.6 86.6
LLC 94.0 97.5 95.7 92.7 97.2 94.9 87.3 86.0 86.6 87.9 70.5 78.2 90.7 87.8 89.2
RLC 98.4 98.8 98.6 97.5 97.2 97.4 98.0 82.6 89.6 97.0 71.6 82.3 97.8 87.5 92.4
Macro avg. 96.3 96.2 96.2 95.5 95.4 95.4 88.1 87.0 87.1 82.9 78.5 78.8 90.0 89.3 89.4
IV-A Dataset Processing
To evaluate the performance of the model proposed in this paper, the highD dataset was employed for both training and testing. The highD dataset is a large-scale natural vehicle trajectory dataset collected on German highways. It encompasses 16.5 hours of data from six distinct locations, involving 110,000 vehicles, covering a total distance of 45,000 km, and capturing 5,600 documented complete lane changes. The highD dataset comprises data extracted from 60 recordings. For our study, we select data from the first 50 recordings for training the model, while the remaining 10 recordings are reserved for testing. Furthermore, we adopt a target-centric strategy that normalizes all position inputs to the coordinate system centered on the current frame position of the target vehicle.
In our work, the lane change frame $\textit{T}_{lc}$ is defined as the frame in which the $lane\_id$ changes. The advanced prediction time T denotes the temporal interval between the lane change frame and the current frame, and it can be expressed as $\textit{T}=\textit{T}_{lc}-\textit{T}_{current}$, as illustrated in the Fig. 2. For the purpose of training and evaluating our proposed LC-LLM model, we extract both lane change (LC) and lane keeping (LK) scenarios from the highD dataset. The LK scenarios comprise data samples wherein the $lane\_id$ remains constant throughout, while the LC scenarios consist of data samples where the advanced prediction time T falls within the range of $[0,4]$ seconds. Furthermore, in order to evaluate the performance of our LC-LLM in different T, we divide LC scenarios into four parts according to the four intervals of $\textit{T}(\textit{T}\in[0,1],\,\textit{T}\in(1,2],\,\textit{T}\in(2,3],\,\textit{T}\in(3,4])$. For the training dataset, we randomly select 48000 lane keeping (LK) samples, 12000 left lane change (LLC) samples from each T interval, and 12000 right lane change (RLC) samples from each T interval, resulting in a total of 144000 samples. Similarly, for the testing dataset, we randomly select 8000 LK samples, 2000 LLC samples from each T interval, and 2000 RLC samples from each T interval, accumulating a total of 24000 samples.
IV-B Experimental Setup
We utilize the training dataset extracted from the highD dataset to train our proposed LC-LLM model, as well as the reimplemented baseline models: an LSTM-based model[22] and a Transformer-based model[3]. All the models are implemented using the PyTorch framework[39]. Additionally, we employ the DeepSpeed [40] library for distributed training of our LC-LLM model. The training process is conducted on eight A800 GPUs and takes 8 hours for the 7B model. The hyperparameters used for model training are detailed in Table I. To ensure the stability of the experimental results, each experiment is repeated five times, and the average value is reported as the final result.
(a) Lateral RMSE $\downarrow$
(b) Longitudinal RMSE $\downarrow$
Figure 6: Comparison of the trajectory prediction performance of the proposed LC-LLM model and baseline models on the highD dataset.
Figure 7: Prompt ablation studies. The ‘Base’ refers to the model results obtained using only the target vehicle information for prompting. The ‘+Map info’, ‘+SV info’, and ‘+CoT’ indicate the model results when map information, surrounding vehicle information, and CoT statements are sequentially incorporated, respectively.
IV-C Evaluation Metrics
Similar to previous works, we utilize precision, recall, F1 score, and macro average metrics to evaluate the performance of the intention prediction task. Additionally, we use the root mean square error (RMSE) metric to evaluate the performance of the trajectory prediction task.
IV-C1 Evaluation Metrics of Intention Prediction
  • Precision: The ratio of correctly predicted positive instances to the total predicted positives. It is calculated as
    Precision=True PositivesTrue Positives + False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives + False Positives}}.$
  • Recall: The ratio of correctly predicted positive instances to the total actual positives. It is calculated as
    Recall=True PositivesTrue Positives + False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives + False Negatives}}.$
  • F1 Score: The harmonic mean of precision and recall. It is calculated as
[…]
  • Macro Avg: The average precision, recall, and F1 score calculated across all classes. It is calculated as the arithmetic mean of precision, recall, and F1 score.
IV-C2 Evaluation Metrics of Trajectory Prediction
  • RMSE (Root Mean Squared Error): A measure of the average deviation between predicted and actual values. It is calculated as
    $\text{RMSE}=\sqrt{\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{n}},$
    where $y_{i}$ is the actual value, $\hat{y}_{i}$ is the predicted value, and $n$ is the number of instances.
In trajectory prediction evaluation, we use RMSE (lat) to assess lateral prediction error and RMSE (lon) to assess longitudinal prediction error.
IV-D Overall Performance
Figure 8: Some visual examples demonstrating the interpretability of our LC-LLM model. In each example, the image above depicts the driving scenario, and the text below presents the corresponding output of our model, which includes the prediction results and their explanations.
We evaluate our proposed LC-LLM model through two tasks, one is intention prediction and the other is trajectory prediction. The results are as follows.
IV-D1 Intention Prediction Results and Analysis
Table II shows the overall performance of the proposed LC-LLM model as well as baseline models in intention prediction with four different advanced prediction time T. We can note that our proposed LC-LLM significantly outperforms baseline models for all T, especially for the longer T. For example, when $\textit{T}\in(2,3]$, the average F1 score of LC-LLM-7B is 2.1% higher than that of the Transformer model and 3.6% higher than that of the LSTM. Subsequently, when considering $\textit{T}\in(3,4]$, the average F1 score of LC-LLM-7B surpasses that of the Transformer model by 1.3% and outperforms the LSTM model by 6.3%. Furthermore, in comparison to the LC-LLM-7B model, the LC-LLM-13B model demonstrates improvements across all advanced prediction times. However, upon scaling the model parameters of LC-LLM to 70B, a decline in intention prediction performance for the longer T is observed. The decrease in performance may be due to not having enough data for fine-tuning. The scaling laws[41] suggest that as the number of parameters in a model increases, it should be accompanied by a corresponding increase in the amount of training data for optimal performance. In this case, the LC-LLM with 70B parameters might not have had sufficient fine-tuning data, leading to suboptimal results as the model couldn’t fully utilize its expanded capacity.
TABLE III: Multi-task Ablation Study.
Task Intention result Trajectory result
F1 (avg) RMSE(lat) RMSE(lon)
Only Intention 0.861
Only Trajectory 0.542 2.032
Intention+Trajectory 0.886 0.478 1.198
TABLE IV: Zero-shot Study.
Method Intention Prediction Trajectory Prediction
F1 score(avg) Failed cases Lateral RMSE(4s) Longitudinal RMSE(4s) Failed cases
Zero-shot(Llama-7B-chat) 0.218 457 2.531 53.795 2928
Fine-tune(LC-LLM-7B) 0.886 0 0.478 1.198 0
IV-D2 Trajectory Prediction Results and Analysis
We assessed the performance of our LC-LLM models in the trajectory prediction task and compared them with baseline models. The results are depicted in Fig. 6. The left panel illustrates the trends of lateral RMSE for trajectory points across different models and prediction horizons, while the right panel shows the trends of longitudinal RMSE. The results indicate a notable improvement in both lateral and longitudinal RMSE for our LC-LLM models in comparison to the baselines. Notably, with larger prediction horizons, our model demonstrates increasing advantages over the baselines in longitudinal RMSE, indicating better robustness concerning long-term prediction horizons. Furthermore, the observed trend suggests that the performance of our LC-LLM model improves as its number of parameters grows, which is particularly evident in the right panel. The relatively small improvement in the left panel is mainly attributed to the fact that the lateral vehicle deviation is minimal in the highway driving scenario, limiting the model’s ability to demonstrate performance differences.
The results of lane change intention prediction task and trajectory prediction task demonstrate that our LC-LLM outperforms all baseline models in long-term (4 s) prediction, achieving a 3.1% improvement in lane change intention prediction, a 19.4% improvement in lateral trajectory prediction, and a 38.1% improvement in longitudinal trajectory prediction, respectively.
IV-E Ablation Study
IV-E1 Prompt Ablation Study
We explored the impact of incorporating contextual prompts on the intention prediction task and the trajectory prediction task, and the results are illustrated in Fig. 7. A series of ablation experiments were conducted, progressively introducing different sources of information—Target Vehicle Information, Map Information, Surrounding Vehicle Information, and Chain-of-Thought statements. The comprehensive model, integrating all contextual prompts, exhibited superior performance compared to the base model with only Target Vehicle Information, achieving an impressive 19.08% improvement in F1 score, a substantial 22.65% reduction in lateral RMSE, and a remarkable 67.85% decrease in longitudinal RMSE. Specifically, the addition of Map Information significantly reduces the lateral error in trajectory prediction, while the inclusion of Surrounding Vehicle Information proved crucial for both intention prediction and future trajectory forecasting. Furthermore, the incorporation of Chain-of-Thought statements yielded notable enhancements in both lane change intention prediction and trajectory prediction. These findings suggest the importance of considering diverse contextual information in prompt engineering with LLM.
IV-E2 Multi-task Ablation Study
We further examined the efficacy of multi-task learning in concurrently predicting lane change intentions and future trajectories, as depicted in Table III. The findings demonstrate superior performance in the multi-task setting compared to the isolated trajectory prediction task, emphasizing the benefits of integrating high-level intention prediction for achieving more accurate trajectory prediction. Furthermore, in comparison to the standalone intention prediction task, adding trajectory prediction not only elevates the overall task complexity but also enables the model to capture more nuanced details, consequently, enhancing the accuracy of intention prediction tasks.
IV-F Interpretability Study
In order to demonstrate the interpretability of our LC-LLM, we visualize several sample scenarios alongside the corresponding outputs from our model, as depicted in Fig. 8. The visualization illustrates that our LC-LLM is capable of forecasting lane change intentions and future trajectories of the target vehicle while also providing explanations for the prediction results. Additionally, the analysis of our LC-LLM’s output indicates that our model not only focuses on the state information of the target vehicle itself but also considers the interactive information of surrounding vehicles to predict the target vehicle’s lane change intention and future trajectory. For example, as shown in the second case in the figure, the target vehicle exhibits a low velocity with no vehicles ahead and no apparent interaction with adjacent vehicles. The explanation given by our LC-LLM is consistent with this driving scenario. In the third case, our LC-LLM elucidates that given the high velocity of the left-front vehicle and the absence of other vehicles on the left, the target vehicle has an opportunity to execute a left lane change. Moreover, in the fourth case, with a slow-moving truck positioned ahead, our LC-LLM predicts that the target vehicle may make a left lane change to avoid collision with the truck. The explanations provided by our model for these scenarios serve as evidence that our LC-LLM possesses a deep understanding of the driving scenario and makes reasonable predictions for the target vehicle. In contrast to previous methodologies that solely generate prediction results, our LC-LLM model not only produces prediction results but also provides corresponding explanations, thereby exhibiting better interpretability than previous methods.
IV-G Zero-shot Study
Table IV presents a comparative analysis between zero-shot and fine-tuned experiments across two tasks. The fine-tuned model demonstrated substantial improvements, yielding markedly higher F1 scores and significantly lower RMSE compared to the zero-shot experiment. These results underscore the efficacy of model fine-tuning in injecting domain-specific knowledge into Large Language Models. This affirms the suitability of fine-tuning as the preferred approach for enabling Large Language Models (LLMs) to excel in specific domains in future applications.
V Conclusion
In this paper, we introduce LC-LLM, an explainable lane change prediction model that not only forecasts lane change intentions and trajectories but also provides explanations for its predictions. We reconceptualize the task of lane change prediction as a language modeling problem and employ a supervised fine-tuning technique to fine-tune the LLM specifically for this task. In this way, we successfully leverage the strong common sense reasoning capabilities and self-explanation abilities of the LLM to address the challenge of lane change prediction. Our extensive experiments conducted on highD dataset show that our LC-LLM improves the accuracy of predicting lane change intention and trajectory as well as significantly augments the interpretability of the prediction results.
Currently, this paper only verifies the effectiveness of our approach in highway scenarios. In the future, our approach should be extended to a wider range of traffic scenarios, and additional experiments on urban scene datasets should be conducted. Furthermore, our LC-LLM exhibit longer inference time compared to baseline models, making the compression of LLMs through knowledge distillation an important area for future research. Moreover, the proposed approach adopts a target-centric strategy to make predictions for one interested vehicle, which results in considerable computational demands when there are multiple interested vehicles in the same scene. Therefore, developing a method that can simultaneously forecast lane change intentions and trajectories for multiple vehicles is an essential direction for future work.
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grant 52302379, Guangzhou Basic and Applied Basic Research Projects under Grants 2023A03J0106 and 2024A04J4290, Guangdong Province General Universities Youth Innovative Talents Project under Grant 2023KQNCX100, and Guangzhou Municipal Science and Technology Project 2023A03J0011.